# Gradient Descent for Noisy Optimization

### v4

@@ 4.1 Example of Proposition 1
Figure 1 shows, for $\delta=0.51$ and for $\delta=0.95$, the evolution of $\mathbb{E}[\|\hat{\theta}_{J,B}-\theta_{\star}\|]$ and of $\mathbb{E}[J_{B}]$ as $B$ increases. For the two considered values of $\delta$ the observed rate at which $\mathbb{E}[\|\hat{\theta}_{J,B}-\theta_{\star}\|]$ decreases to zero is indistinguishable from the optimal $B^{-1/2}$ convergence rate. In addition, in this experiment the value of $\delta$ appears to have little impact on $\mathbb{E}[\|\hat{\theta}_{J,B}-\theta_{\star}\|]$. As expected, we see in Figure 1 that the number of calls to Algorithm 3 is much larger for $\delta=0.95$ than for $\delta=0.51$, with a value of $\mathbb{E}[J_{B}]$ ranging from 20.48 to 34.04 for the former choice of $\delta$ and ranging from 2.37 to 3.28 for the latter. Visually, we observe (at least for $\delta=0.95$) that $\mathbb{E}[J_{B}]$ grows with $B$ as $\log B$. The coefficient of correlation between the variables $(\mathbb{E}[J_{B}])_{B\geq 1}$ and $(\log B)_{B\geq 1}$ confirms this observation, since for this experiment we have $\rho_{J_{B}}=0.99$ when $\delta=0.95$ and $\rho_{J_{B}}=0.78$ when $\delta=0.51$.
@@ 4.2 A multivariate extension of the example of Proposition 1
$X\sim\mathrm{Poisson}(1),\quad W|X\sim\mathcal{U}([-1,1]^{d-1}),\quad Y|(W,X)\sim\mathrm{Poisson}(a^{\top}_{\star}W)$
@@
$f(\theta,z)=-z_{2}\theta^{\top}z_{1}+\exp(\theta^{\top}z_{1}),\quad\theta\in\mathbb{R}^{d},\quad z\in\mathsf{Z}:=\mathbb{R}^{d+1}.$
It is easily checked that the resulting objective function $F$ is such that A1-A2 hold with $\alpha=2$ and that $\theta_{\star}=(0,a_{\star})$. Compared to the example of Proposition 1, we expect that by increasing the dimension of the search space we make the optimization problem more challenging and thus that the conclusions of Proposition 1-2 hold for this example. For this experiment the value $a_{\star}$ is a random draw from the $\mathcal{N}_{d-1}(0,\mathbbm{1}_{d-1})$ distribution and, as in the previous subsection, the results are presented for $(\alpha^{\prime},\delta)=(1,0.51)$ and for $(\alpha^{\prime},\delta)=(1,0.95)$. Finally, Algorithm 4 is implemented using $\theta_{0}=(1,\dots,1)$ as starting value.
From Figure 2, which shows the evolution of $\mathbb{E}[\|\hat{\theta}_{J,B}-\theta_{\star}\|]$ and of $\mathbb{E}[J_{B}]$ as $B$ increases, we see that the behaviour of these two quantities is similar to what we observed in the previous subsection. The main difference is however that, because of the larger dimension of $\theta$, the $B^{-1/2}$ convergence rate for $\hat{\theta}_{J,B}$ is observed only for $B$ large enough (roughly for $B>10^{6.5}$). The study of the dependence of the performance of Algorithm 4 to the dimension $d$ of $\theta$ is left for future research. Finally, in Figure 2 we also observe that $\mathbb{E}[J_{B}]$ appears to increase with $B$ as $\log B$, and indeed, for this example, the coefficient of correlation between $(\mathbb{E}[J_{B}])_{B\geq 1}$ and $(\log B)_{B\geq 1}$ is $\rho_{J_{B}}=0.98$ for $\delta=0.95$ and $\rho_{J_{B}}=0.70$ for $\delta=0.51$.
@@ 4.3 Example of Proposition 1 with an extra heavy tailed noise
The plots in the first two columns of Figure 3 are obtained for $(\alpha^{\prime},\delta)=(0.5,0.41)$ and for $(\alpha^{\prime},\delta)=(0.5,0.95)$. From these plots we see that the convergence rate of $\hat{\theta}_{J,B}$ is undistinguishable from the $B^{-\frac{\alpha^{\prime}}{1+\alpha^{\prime}}}=B^{-1/3}$ convergence rate, confirming the result of Theorem 2. As for the examples of Sections 4.1-4.2, the choice of $\delta$ appears to have only a limited impact on the estimation error. We also observe that for the two considered values of $\delta$ the quantity $\mathbb{E}[J_{B}]$ seems to grow with $B$ as $\log B$, an observation which is confirmed by the fact that $\rho_{J_{B}}=0.85$ for $\delta=0.41$ and $\rho_{J_{B}}=0.98$ for $\delta=0.95$. The plots in last column of Figure 3 study the behaviour of $\hat{\theta}_{J,B}$ when $(\alpha^{\prime},\delta)=(1,0.95)$, in which case A2 does not hold for $\alpha=\alpha^{\prime}$. If $\hat{\theta}_{J,B}$ appears to roughly converge to $\theta_{\star}$ at speed $B^{-1/3}$, we observe that the convergence becomes slower for values of $B$ greater than $10^{6.7}$. Remark that, from the discussion of Section 3.2.2, it was expected that the convergence rate of $\hat{\theta}_{J,B}$ is slower when A2 does not hold for the chosen value of $\alpha^{\prime}$ than when it does. Finally, for this value of $(\alpha^{\prime},\delta)$ the quantity $\mathbb{E}[J_{B}]$ also seems to grow as $\log B$, an observation which is confirmed by the large value $\rho_{J_{B}}=0.98$ obtained for the coefficient of correlation between $(\mathbb{E}[J_{B}])_{B\geq 1}$ and $(\log B)_{B\geq 1}$.
@@ 5 Concluding remarks
The proposed algorithm (Algorithm 4) requires that GD-BLS can be called at most $J$ times, with $J$ a user-specified parameter. We conjecture that this condition can be removed and, assuming that A2 holds for $\alpha=\alpha^{\prime}$, that the resulting estimator $\hat{\theta}_{B}$ of the global minimiser $\theta_{\star}$ converges at speed $B^{-\frac{\alpha^{\prime}}{1+\alpha^{\prime}}(1-\delta^{j}_{B})}$ for some sequence $(j_{B})_{B\geq 1}$ such that $\lim_{B\rightarrow\infty}j_{B}=\infty$. Based on our numerical experiments, we predict that $j_{B}=a+b\log B$ for some constants $a\in[0,\infty)$ and $b\in(0,\infty)$ depending on the choice of $\delta$. Confirming these results theoretically is left for future research.
@@ A.1 Proof of Proposition 2
$\mathcal{L}_{\alpha,\epsilon}(\theta)\geq K_{\frac{1}{2}r_{\alpha,\epsilon,\theta},\alpha}(\theta),\quad\forall\theta\geq c_{\alpha,\epsilon}^{\prime}$ (11)
@@ A.2.2 Proof of Lemma 1
Since, as shown above, $\mathbb{E}[M^{(1)}_{\Theta}(Z)]<\infty$ while, by assumption, $\mathbb{E}[\|\nabla f(\theta_{\star},Z)\|]<\infty$, it follow that $\mathbb{E}[M^{(2)}_{\Theta}(Z)]<\infty$.
The result of the lemma follows. ∎
@@ A.2.3 Proof of Lemma 2
where the first inequity holds by Jensen’s inequality and the second one holds under A2. Since $\mathbb{E}[M_{\Theta}(Z)]<\infty$, this shows that the mapping $\nabla^{2}F:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d\times d}$ is locally Lipschitz and thus continuous. The proof of the lemma is complete. ∎
@@ A.2.6 Proof of Lemma 5
Remark first that under the assumptions of the lemma we have $\mathbb{E}[\nabla f(\theta_{\star},Z)]=0$ by Lemma 2. Hence, since $\mathbb{E}[\|\nabla f(\theta_{\star},Z)\|^{1+\alpha}]<\infty$ by assumption, it follows from Bubeck et al., (2013, Lemma 3) that there exists a constant $C^{*}<\infty$ such that
@@ A.2.7 Proof of Proposition 4
We consider the optimization problem (1) with $d=1$, with $\mathsf{Z}=\mathbb{R}$, with $f:\mathbb{R}\times\mathsf{Z}\rightarrow\mathbb{R}$ defined by $f(\theta,z)=\theta^{2}-2\theta z$, $(\theta,z)\in\mathbb{R}^{2}$ and with $Z\sim P$ for some probability distribution $P$. Then, it is direct to check that for any distribution $P$ such that $\mathbb{E}[|Z|^{1+\alpha^{\prime}}]<\infty$ Assumptions A1-A2 hold with $\alpha=\alpha^{\prime}$, and that we have $\theta_{\star}=\mathbb{E}[Z]$. For all $B\in\mathbb{N}$ let $n_{B}$ denote the number of distinct elements of $(Z_{i})_{i\geq 1}$ used to define $\hat{\theta}_{B}$. Then, by Devroye et al., (2016, Theorem 3.1), there exists a distribution $P$ of $Z$ such that $\mathbb{E}[|Z|^{1+\alpha^{\prime}}]<\infty$ and such that, for some constants $\delta\in(0,1)$ and $C\in(0,\infty)$, we have
$\limsup_{B\rightarrow\infty}\mathbb{P}\Big{(}|\hat{\theta}_{B}-\mathbb{E}[Z]|>Cn_{B}^{-\frac{\alpha}{1+\alpha}}\big{)}\geq\delta.$
The result of the proposition follows upon noting that $n_{B}\leq B$ for all $B\in\mathbb{N}$. ∎
@@ A.2.8 Proof of Proposition 5
Since by assumption $g_{\star}>-\infty$ and $g$ is convex it trivially exists an $x_{\star}\in\mathbb{R}^{d}$ such that $g_{\star}=g(x_{\star})$.
Next, remark that since $g$ is convex for all $t\geq 0$ we have
$g(x_{\star})\geq g(x_{t})+\nabla g(x_{t})^{\top}(x_{\star}-x_{t})$
@@
In addition, remark that
$\|x_{t+1}-x_{\star}\|^{2}=\|x_{t}-x_{\star}\|^{2}+v_{t}^{2}\|\nabla g(x_{t})\|^{2}-2v_{t}\nabla g(x_{t})^{\top}(x_{t}-x_{\star}),\quad\forall t\geq 0.$ (23)
Let $t\geq 0$ and note that
@@
where the first inequality uses (2), the second inequality uses (22) and the first equality uses (23).
@@ A.2.9 Proof of Proposition 6
For a given $x_{0}\in\mathbb{R}^{d}$ the result of Proposition 6 follows by applying the following proposition with $\tilde{K}=\{x\in\mathbb{R}^{d}:\,g(x)\leq g(x_{0})\}$, noting that this set is indeed compact and convex since $g$ is assumed to be a convex function.
Proposition 7.
Let $g:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be a convex, differentiable and locally $L$-smooth function such that $g_{\star}>-\infty$. Let $x_{0}\in\mathbb{R}^{d}$, $\beta\in(0,1)$ and, for all $T\in\mathbb{N}$, let $x_{T}$ be as computed by Algorithm 1 where Algorithm 2 is used to compute $v_{t}$ for all $t\geq 1$. In addition, let $\tilde{K}\subset\mathbb{R}^{d}$ be a compact and convex set such that $x_{T}\in\tilde{K}$ for all $T\geq 1$,
$K=\Big{\{}x\in\mathbb{R}^{d}:\,\exists x^{\prime}\in\tilde{K}\text{ such that }\|x-x^{\prime}\|\leq\sup_{\tilde{x}\in\tilde{K}}\|\nabla g(\tilde{x})\|\Big{\}}$
and, noting that $K$ is a compact set, let $L_{K}\in[1,\infty)$ be such that $\|\nabla g(x)-\nabla g(x^{\prime})\|\leq L_{K}\|x-x^{\prime}\|$ for all $x,x^{\prime}\in K$. Then, (2) holds with $c=\beta/L_{K}$.
@@
Let $K$ and $\tilde{K}$ be as in the statement of the proposition.
Remark first that, since $g$ is $L_{K}$-smooth on the convex set $K$, we have
$g(y)\leq g(x)+\nabla g(x)^{\top}(y-x)+\frac{L_{K}}{2}\|y-x\|^{2},\quad\forall x,y\in K$ (26)
and remark that $x-v\nabla g(x)\in K$ for all $x\in\tilde{K}$ and $v\in(0,1]$. We now let $x\in\tilde{K}\subset K$ and $v\in(0,1]$. Then, since by (26) we have
@@
then $v\geq\beta/L_{K}$. Since by assumption $x_{T}\in\tilde{K}$ for all $T\geq 0$, this shows that $v_{T}\geq\beta/L_{K}$ for all $T\geq 1$ and the proof of the proposition is complete. ∎
@@ A.3.2 A first key property of Algorithm 3
Assume that A2 holds. Let $\beta\in(0,1)$, $(n_{B})_{B\geq 1}$ be a sequence in $\mathbb{N}$, $(\tau_{B})_{B\geq 1}$ be a sequence in $[0,\infty)$ and $(\tilde{B}_{B})_{B\geq 1}$ be a sequence of $\mathbb{N}_{0}$-valued random variables. In addition, for all $B\in\mathbb{N}$ and $\theta_{0}\in\mathbb{R}^{d}$, let
@@
and denote by $\{\theta_{\theta_{0},B,t}\}_{t=0}^{T_{\theta_{0},B}}$ the sequence in $\mathbb{R}^{d}$ generated by Algorithm 3 to compute $\hat{\theta}_{\theta_{0},B}$ from $\theta_{0}$ and (which is therefore such that $\theta_{\theta_{0},B,0}=\theta_{0}$ and such that $\theta_{\theta_{0},B,T_{\theta_{0},B}}=\hat{\theta}_{\theta_{0},B}$). Finally, let $x\in\mathbb{R}^{d}$ and, for all $j\in\mathbb{N}$, let
$\Theta_{j,x}=\big{\{}\theta\in\mathbb{R}^{d}:\,F(\theta)\leq F(x)+j\big{\}}.$
Then, for all $j\in\mathbb{N}$, there exists a sequence $(\Omega^{(1)}_{j,B})_{B\geq 1}$ in $\mathcal{F}$ such that $\lim_{B\rightarrow\infty}\mathbb{P}(\Omega^{(1)}_{j,B})=1$ and such that
@@
Let $j\in\mathbb{N}$,
$\widetilde{\Theta}_{j,x}=\Big{\{}\theta\in\mathbb{R}^{d}:\,\exists\theta^{\prime}\in\Theta_{j+1,x}\text{ such that }\|\theta-\theta^{\prime}\|\leq\sup_{\tilde{\theta}\in\Theta_{j+1,x}}\|\nabla F(\tilde{\theta})\|+1\Big{\}}$
@@
$\Omega^{(1)}_{j,B}=\Big{\{}\omega\in\Omega:,\max_{k\in\{0,1\}}\sup_{\theta\in\tilde{\Theta}_{j,x}}\|\nabla^{k}F^{\omega}_{n_{B}}(\theta)-\nabla^{k}F(\theta)\|\leq 1/2\Big{\}}.$
By Lemma 2, the function $F$ is continuous on $\mathbb{R}^{d}$ and thus the set $\Theta_{j+1,x}$ is compact. Therefore, using again Lemma 2, we have $\sup_{\tilde{\theta}\in\Theta_{j+1,x}}\|\nabla F(\tilde{\theta})\|<\infty$ and thus the set $\widetilde{\Theta}_{j,x}$ is also compact. Consequently, by Lemma 3, we have $\lim_{B\rightarrow\infty}\mathbb{P}(\Omega^{(1)}_{j,B})=1$.

### v5

@@ 4.1 Example of Proposition 1
Figure 1 shows, for $\delta=0.51$ and for $\delta=0.95$, the evolution of $\mathbb{E}[\|\hat{\theta}_{J,B}-\theta_{\star}\|]$ and of $\mathbb{E}[J_{B}]$ as $B$ increases. For the two considered values of $\delta$ the observed rate at which $\mathbb{E}[\|\hat{\theta}_{J,B}-\theta_{\star}\|]$ decreases to zero is indistinguishable from the optimal $B^{-1/2}$ convergence rate. In addition, in this experiment the value of $\delta$ appears to have little impact on $\mathbb{E}[\|\hat{\theta}_{J,B}-\theta_{\star}\|]$. As expected, we see in Figure 1 that the number of calls to Algorithm 3 is much larger for $\delta=0.95$ than for $\delta=0.51$, with a value of $\mathbb{E}[J_{B}]$ ranging from 20.48 to 34.04 for the former choice of $\delta$ and ranging from 2.37 to 3.28 for the latter. Visually, we observe (at least for $\delta=0.95$) that $\mathbb{E}[J_{B}]$ grows with $B$ as $\log(B)$. The coefficient of correlation between the variables $(\mathbb{E}[J_{B}])_{B\geq 1}$ and $(\log(B))_{B\geq 1}$ confirms this observation, since for this experiment we have $\rho_{J_{B}}=0.99$ when $\delta=0.95$ and $\rho_{J_{B}}=0.78$ when $\delta=0.51$.
@@ 4.2 A multivariate extension of the example of Proposition 1
$W\sim\mathrm{Poisson}(1),\quad X|W\sim\mathcal{U}([-1,1]^{d-1}),\quad Y|(W,X)\sim\mathrm{Poisson}(a^{\top}_{\star}X)$
@@
$f(\theta,z)=-z_{2}\theta^{\top}z_{1}+\exp(\theta^{\top}z_{1}),\quad\theta\in\mathbb{R}^{d},\quad z_{1}\in\mathbb{R}^{d},\quad z_{2}\in\mathbb{R},\quad z=(z_{1},z_{2})\in\mathsf{Z}:=\mathbb{R}^{d+1}.$
It is easily checked that the resulting objective function $F$ is such that A1-A2 hold with $\alpha=1$ and that $\theta_{\star}=(0,a_{\star})$. Compared to the example of Proposition 1, we expect that by increasing the dimension of the search space we make the optimization problem more challenging and thus that the conclusions of Proposition 1-2 hold for this example. For this experiment the value $a_{\star}$ is a random draw from the $\mathcal{N}_{d-1}(0,\mathbbm{1}_{d-1})$ distribution and, as in the previous subsection, the results are presented for $(\alpha^{\prime},\delta)=(1,0.51)$ and for $(\alpha^{\prime},\delta)=(1,0.95)$. Finally, Algorithm 4 is implemented using $\theta_{0}=(1,\dots,1)$ as starting value.
From Figure 2, which shows the evolution of $\mathbb{E}[\|\hat{\theta}_{J,B}-\theta_{\star}\|]$ and of $\mathbb{E}[J_{B}]$ as $B$ increases, we see that the behaviour of these two quantities is similar to what we observed in the previous subsection. The main difference is however that, because of the larger dimension of $\theta$, the $B^{-1/2}$ convergence rate for $\hat{\theta}_{J,B}$ is observed only for $B$ large enough (roughly for $B>10^{6.5}$). The study of the dependence of the performance of Algorithm 4 to the dimension $d$ of $\theta$ is left for future research. Finally, in Figure 2 we also observe that $\mathbb{E}[J_{B}]$ appears to increase with $B$ as $\log(B)$, and indeed, for this example, the coefficient of correlation between $(\mathbb{E}[J_{B}])_{B\geq 1}$ and $(\log(B))_{B\geq 1}$ is $\rho_{J_{B}}=0.98$ for $\delta=0.95$ and $\rho_{J_{B}}=0.70$ for $\delta=0.51$.
@@ 4.3 Example of Proposition 1 with an extra heavy tailed noise
The plots in the first two columns of Figure 3 are obtained for $(\alpha^{\prime},\delta)=(0.5,0.41)$ and for $(\alpha^{\prime},\delta)=(0.5,0.95)$. From these plots we see that the convergence rate of $\hat{\theta}_{J,B}$ is undistinguishable from the $B^{-\frac{\alpha^{\prime}}{1+\alpha^{\prime}}}=B^{-1/3}$ convergence rate, confirming the result of Theorem 2. As for the examples of Sections 4.1-4.2, the choice of $\delta$ appears to have only a limited impact on the estimation error. We also observe that for the two considered values of $\delta$ the quantity $\mathbb{E}[J_{B}]$ seems to grow with $B$ as $\log(B)$, an observation which is confirmed by the fact that $\rho_{J_{B}}=0.85$ for $\delta=0.41$ and $\rho_{J_{B}}=0.98$ for $\delta=0.95$. The plots in last column of Figure 3 study the behaviour of $\hat{\theta}_{J,B}$ when $(\alpha^{\prime},\delta)=(1,0.95)$, in which case A2 does not hold for $\alpha=\alpha^{\prime}$. If $\hat{\theta}_{J,B}$ appears to roughly converge to $\theta_{\star}$ at speed $B^{-1/3}$, we observe that the convergence becomes slower for values of $B$ greater than $10^{6.7}$. Remark that, from the discussion of Section 3.2.2, it was expected that the convergence rate of $\hat{\theta}_{J,B}$ is slower when A2 does not hold for the chosen value of $\alpha^{\prime}$ than when it does. Finally, for this value of $(\alpha^{\prime},\delta)$ the quantity $\mathbb{E}[J_{B}]$ also seems to grow as $\log(B)$, an observation which is confirmed by the large value $\rho_{J_{B}}=0.98$ obtained for the coefficient of correlation between $(\mathbb{E}[J_{B}])_{B\geq 1}$ and $(\log(B))_{B\geq 1}$.
@@ 5 Concluding remarks
The proposed algorithm (Algorithm 4) requires that GD-BLS can be called at most $J$ times, with $J$ a user-specified parameter. We conjecture that this condition can be removed and, assuming that A2 holds for $\alpha=\alpha^{\prime}$, that the resulting estimator $\hat{\theta}_{B}$ of the global minimiser $\theta_{\star}$ converges at speed $B^{-\frac{\alpha^{\prime}}{1+\alpha^{\prime}}(1-j_{B})}$ for some sequence $(j_{B})_{B\geq 1}$ such that $\lim_{B\rightarrow\infty}j_{B}=\infty$. Based on our numerical experiments, we predict that $j_{B}=a+b\log(B)$ for some constants $a\in[0,\infty)$ and $b\in(0,\infty)$ depending on the choice of $\delta$. Confirming these results theoretically is left for future research.
@@ A.1 Proof of Proposition 2
$\mathcal{L}_{\alpha,\epsilon}(\theta)\geq K_{\frac{1}{2}r_{\alpha,\epsilon,\theta},\alpha}(\theta),\quad\forall\theta\geq c_{\alpha,\epsilon}^{\prime}.$ (11)
@@ A.2.2 Proof of Lemma 1
Since, as shown above, $\mathbb{E}[M^{(1)}_{\Theta}(Z)]<\infty$ while, by assumption, $\mathbb{E}[\|\nabla f(\theta_{\star},Z)\|]<\infty$, it follow that $\mathbb{E}[M^{(2)}_{\Theta}(Z)]<\infty$. The result of the lemma follows. ∎
@@ A.2.3 Proof of Lemma 2
where the first inequality holds by Jensen’s inequality and the second one holds under A2. Since $\mathbb{E}[M_{\Theta}(Z)]<\infty$, this shows that the mapping $\nabla^{2}F:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d\times d}$ is locally Lipschitz and thus continuous. The result of the lemma follows. ∎
@@ A.2.6 Proof of Lemma 5
Let $\Theta$ be a compact and convex set such that $\theta_{\star}\in\mathring{\Theta}$ and remark that, by Lemma 2, we have $\mathbb{E}[\nabla f(\theta_{\star},Z)]=0$. Hence, since $\mathbb{E}[\|\nabla f(\theta_{\star},Z)\|^{1+\alpha}]<\infty$ by assumption, it follows from Bubeck et al., (2013, Lemma 3) that there exists a constant $C^{*}<\infty$ such that
@@ A.2.7 Proof of Proposition 4
We consider the optimization problem (1) with $d=1$, with $\mathsf{Z}=\mathbb{R}$, with $f:\mathbb{R}\times\mathsf{Z}\rightarrow\mathbb{R}$ defined by $f(\theta,z)=\theta^{2}-2\theta z$, $(\theta,z)\in\mathbb{R}^{2}$ and with $Z\sim P$ for some probability distribution $P$. Then, it is direct to check that for any distribution $P$ such that $\mathbb{E}[|Z|^{1+\alpha^{\prime}}]<\infty$ Assumptions A1-A2 hold with $\alpha=\alpha^{\prime}$, and that we have $\theta_{\star}=\mathbb{E}[Z]$. For all $B\in\mathbb{N}$ let $\hat{\theta}_{B}$ be an estimator of $\theta_{\star}$ based on $k_{B}\leq B$ distinct elements of $(Z_{i})_{i\geq 1}$. Then, from the proof of Devroye et al., (2016, Theorem 3.1), it is direct to see that there exists a distribution $P$ of $Z$ such that $\mathbb{E}[|Z|^{1+\alpha^{\prime}}]<\infty$ and such that, for some constants $\delta\in(0,1)$ and $C\in(0,\infty)$, we have
$\limsup_{B\rightarrow\infty}\mathbb{P}\Big{(}|\hat{\theta}_{B}-\mathbb{E}[Z]|>Ck_{B}^{-\frac{\alpha}{1+\alpha}}\big{)}\geq\delta.$
The result of the proposition follows upon noting that $k_{B}\leq B$ for all $B\in\mathbb{N}$. ∎
@@ A.2.8 Proof of Proposition 5
Since $g$ is convex on the convex set $K$ and $x_{t}\in K$ for all $t\geq 0$, it follows that
$g(x_{\star})\geq g(x_{t})+\nabla g(x_{t})^{\top}(x_{\star}-x_{t}),\quad\forall t\geq 0$
@@
Remark that
$\|x_{t+1}-x_{\star}\|^{2}=\|x_{t}-x_{\star}\|^{2}+v_{t}^{2}\|\nabla g(x_{t})\|^{2}-2v_{t}\nabla g(x_{t})^{\top}(x_{t}-x_{\star}),\quad\forall t\geq 0$ (23)
and let $t\geq 0$. Then,
@@
where the first inequality uses (2), the second inequality uses (22) and the third equality uses (23).
@@ A.2.9 Proof of Proposition 6
Let $K$ and $\tilde{K}$ be as in the statement of the proposition, and remark that since $g$ is convex and $L_{K}$-smooth on the convex set $K$ we have, by the descent lemma,
$g(y)\leq g(x)+\nabla g(x)^{\top}(y-x)+\frac{L_{K}}{2}\|y-x\|^{2},\quad\forall(x,y)\in K^{2}.$ (26)
Remark also that $x-v\nabla g(x)\in K$ for all $x\in\tilde{K}$ and all $v\in(0,1]$, and let $x\in\tilde{K}\subset K$ and $v\in(0,1]$. Then, since by (26) we have
@@
then $v\geq\beta/L_{K}$. Since $x_{t}\in\tilde{K}$ for all $t\geq 0$, this shows that $v_{t}\geq\beta/L_{K}$ for all $t\geq 1$ and the proof of the proposition is complete. ∎
@@ A.3.2 A first key property of Algorithm 3
Assume that A1-A2 hold. Let $x\in\mathbb{R}^{d}$, $\beta\in(0,1)$, $(n_{B})_{B\geq 1}$ be a sequence in $\mathbb{N}$, $(\tau_{B})_{B\geq 1}$ be a sequence in $[0,\infty)$ and $(\tilde{B}_{B})_{B\geq 1}$ be a sequence of $\mathbb{N}_{0}$-valued random variables. In addition, for all $B\in\mathbb{N}$ and $\theta_{0}\in\mathbb{R}^{d}$, let
@@
and denote by $\{\theta_{\theta_{0},B,t}\}_{t=0}^{T_{\theta_{0},B}}$ the sequence in $\mathbb{R}^{d}$ generated by Algorithm 3 to compute $\hat{\theta}_{\theta_{0},B}$ from $\theta_{0}$ and (which is therefore such that $\theta_{\theta_{0},B,0}=\theta_{0}$ and such that $\theta_{\theta_{0},B,T_{\theta_{0},B}}=\hat{\theta}_{\theta_{0},B}$). Finally, for all $j\in\mathbb{N}_{0}$, let $\Theta_{j,x}=\big{\{}\theta\in\mathbb{R}^{d}:\,F(\theta)\leq F(x)+j\big{\}}$. Then, for all $j\in\mathbb{N}_{0}$, the set $\Theta_{j+1,x}$ is a (non-empty) compact and convex set such that $\theta_{\star}\in\mathring{\Theta}_{j+1,x}$, and there exists a sequence $(\Omega^{(1)}_{j,B})_{B\geq 1}$ in $\mathcal{F}$ such that $\lim_{B\rightarrow\infty}\mathbb{P}(\Omega^{(1)}_{j,B})=1$ and such that
@@
Let $j\in\mathbb{N}_{0}$,
$\widetilde{\Theta}_{j,x}=\Big{\{}\theta\in\mathbb{R}^{d}:\,\exists\theta^{\prime}\in\Theta_{j+1,x}\text{ such that }\|\theta-\theta^{\prime}\|\leq\sup_{\tilde{\theta}\in\Theta_{j+1,x}}\|\nabla F(\tilde{\theta})\|+1\Big{\}}$ (34)
@@
$\Omega^{(1)}_{j,B}=\Big{\{}\omega\in\Omega:\max_{k\in\{0,1\}}\sup_{\theta\in\tilde{\Theta}_{j,x}}\|\nabla^{k}F^{\omega}_{n_{B}}(\theta)-\nabla^{k}F(\theta)\|\leq 1/2\Big{\}}.$
Under A1 the function $F$ is strictly convex on $\mathbb{R}^{d}$ and thus the set $\Theta_{j+1,x}$ is convex and bounded. Therefore, since by Lemma 2 the function $F$ is also continuous on $\mathbb{R}^{d}$, the set $\Theta_{j+1,x}$ is compact and such that $\theta_{\star}\in\mathring{\Theta}_{j+1,x}$.
To show the second part of lemma remark that since $\Theta_{j+1,x}$ is a compact (and convex) set and $F$ is continuous by Lemma 2, it follows that $\sup_{\tilde{\theta}\in\Theta_{j+1,x}}\|\nabla F(\tilde{\theta})\|<\infty$ and thus $\widetilde{\Theta}_{j,x}$ is also a compact (and convex) set. Consequently, by Lemma 3, we have $\lim_{B\rightarrow\infty}\mathbb{P}(\Omega^{(1)}_{j,B})=1$.